OpenStudy (anonymous):

Evaluate the polynomial for  x=2  and for  x=−2. −3x3+8x2−5x−9

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

i gutchu

OpenStudy (anonymous):

ight boi listen up quickly

OpenStudy (anonymous):

x3 + 3x2 + 5x – 4) – (3x3 – 8x2 – 5x + 6) = (x3 + 3x2 + 5x – 4) – 1(3x3 – 8x2 – 5x + 6) = (x3 + 3x2 + 5x – 4) – 1(3x3) – 1 (–8x2) – 1(–5x) – 1(6) = x3 + 3x2 + 5x – 4 – 3x3 + 8x2 + 5x – 6 = x3 – 3x3 + 3x2 + 8x2 + 5x + 5x – 4 – 6 = –2x3 + 11x2 + 10x –10

OpenStudy (anonymous):

2x3 – 2x2 + 19x – 10